find the derivative of the following functions wrt x
question = y = (sinx )^x

Grade:12

1 Answers

Robin Chourasia
10 years ago
y= (sinx)^x

Taking log,

lny = x ln(sinx)

Taking derivative wrt x,

1/y dy/dx = x (1/sinx)cosx + ln(sinx)

Put y and simplyfy.

Thanks and regards,
Robin
askIITian Faculty.